# Scaling Laws with Vocabulary: Larger Models Deserve Larger Vocabularies

### v2

@@ Abstract
Research on scaling large language models (LLMs) has primarily focused on model parameters and training data size, overlooking the role of vocabulary size. We investigate how vocabulary size impacts LLM scaling laws by training models ranging from 33M to 3B parameters on up to 500B characters with various vocabulary configurations. We propose three complementary approaches for predicting the compute-optimal vocabulary size: IsoFLOPs analysis, derivative estimation, and parametric fit of the loss function. Our approaches converge on the same result that the optimal vocabulary size depends on the available compute budget and that larger models deserve larger vocabularies. However, most LLMs use too small vocabulary sizes. For example, we predict that the optimal vocabulary size of Llama2-70B should have been at least 216K, 7 times larger than its vocabulary of 32K. We validate our predictions empirically by training models with 3B parameters across different FLOPs budgets. Adopting our predicted optimal vocabulary size consistently improves downstream performance over commonly used vocabulary sizes. By increasing the vocabulary size from the conventional 32K to 43K, we improve performance on ARC-Challenge from 29.1 to 32.0 with the same 2.3e21 FLOPs. Our work emphasizes the necessity of jointly considering model parameters and vocabulary size for efficient scaling.
@@ 1 Introduction
Large language models (LLMs) achieve remarkable performance by pre-training on vast text corpora using massive computational resources [46]. Extensive prior work on LLMs has focused on deriving so-called scaling laws: a set of empirical formulas to predict how model performance scales, mainly as computing floating-point operations (FLOPs), model parameters, and quantity of training data change [30, 26, 63, 2, 43, 57]. These works show that power-law fits can effectively predict language modeling loss and by extension downstream performance [23, 54]. However, these scaling laws usually disregard the impact of the vocabulary size. For example, in Kaplan et al. [30] only non-vocabulary parameters are considered in their predictive formula. This negligence has resulted in substantial variability in the vocabulary size of current LLMs. For instance, Llama2-7B employs a vocabulary size of 32K [67], while Gemma-7B [64] adopts a much larger vocabulary size of 256K despite both having a similar number of total parameters. This variability in vocabulary sizes across LLMs raises the research question: What is the compute-optimal vocabulary size for an LLM?
The vocabulary size affects performance non-trivially. Intuitively, the optimal vocabulary size should neither be excessively large nor small. A larger vocabulary size improves tokenization fertility, i.e., splitting sentences into fewer tokens, thereby improving the tokenization efficiency. Additionally, a larger vocabulary enhances the representational capacity of the model, enabling it to capture a wider range of concepts and nuances in the corpus. However, the risk of under-fitting representations for rare tokens increases with larger vocabulary sizes, especially in the data-constrained regime [43, 69]. Thus, the optimal vocabulary size needs to be determined by taking the training data and the amount of non-vocabulary parameters into account.
@@ 2.1 Scaling law
Following Radford et al. [50], the loss function is typically the language modeling loss, which can be written as:
@@ 2.2 Scaling law with vocabulary
Attributes
Scaling laws commonly deal with the attributes, model parameters ($N$) and number of training tokens ($D$) [26, 43]. We adapt them for our analysis in the context of vocabulary size. (1) We break down the total model parameters ($N$) into non-vocabulary ($N_{\rm nv}$) and vocabulary parameters ($N_{\rm v}$). To understand the importance of vocabulary parameters, we isolate them from other model parameters, where $N=N_{\rm nv}+N_{\rm v}$. We use $N_{\rm v}=Vd$ to represent both the vocabulary parameters in the output layer  [^1] Vocabulary parameters typically encompass both the word embedding layer and the output layer. In this paper, for clarity and analytical simplicity, we employ $Vd$ rather than $2Vd$ to represent the vocabulary parameters. This methodological choice is predicated on empirical observations: the predominant computational burden, as measured in FLOPs, is associated with the output layer, while the computational cost of the word embedding layer is comparatively insignificant. Consequently, references to vocabulary parameters or their associated FLOPs primarily pertain to those in the output layer, denoted by $Vd$.. Notably, to change $N_{\rm v}$ we only vary the vocabulary size $V$ and take the embedding dimension $d$ as given based on $N_{\rm nv}$ empirically, see §A.5.2 for details. This is based on the observation by Kaplan et al. [30] that the performance of models with varying depth-to-width ratios converges to a single trend. (2) We measure data not in tokens ($D$) but in training characters ($H$). The number of tokens depends on the vocabulary of the tokenizer, thus we need a vocabulary-independent measurement of data. By studying training characters, we can better see how the data volume affects the performance regardless of different vocabulary sizes.
Mapping from training characters ($H$) to tokens ($D$)
As detailed above we measure training data in training characters ($H$). Nonetheless, to connect our findings with existing studies on scaling laws [26, 43], we need to be able to map from $H$ to $D$. This mapping is the tokenizer’s compression ratio which can be computed via $D/H$. The more tokens the tokenizer needs to represent $H$, the larger $D$, and thus it compresses less. We develop a simple function $f(V)$ to estimate this ratio solely from the chosen vocabulary size, $V$. Specifically, we find that a quadratic function on the logarithmic value of $V$ achieves accurate predictions:
@@
Vocabulary-insensitive loss
To fairly assess models that vary in $V$, the commonly used language model loss in Equation 2 is inappropriate. Models trained with larger $V$ naturally have a higher loss, as there are more possibilities in the vocabulary to predict. However, this does not mean that the model is worse. Thus, we need to normalize the loss with respect to the vocabulary size. We reformulate the unigram-normalized metric [53] as a loss function. Suppose we have a $T$-length sequence $w_{1:T}$, we design the unigram-normalized language model loss as:
@@
where $p(w_{i}|V)$ is the frequency of word $w_{i}$ in the tokenized corpus, given the tokenizer with vocabulary size $V$. The loss indicates the improvement in probability that a context-aware language model offers over a unigram model without context, allowing us to assess the language model’s efficacy. Based on theory from prior work [53], the normalized loss $\mathcal{L}_{u}$ remains consistent for a given model with a fixed non-vocabulary component across different vocabulary sizes. The difference of $\mathcal{L}_{u}$ comes from the ability of the language model itself. Compared with $\mathcal{L}$, the value of $\mathcal{L}_{u}$ is much smaller and can be negative as $\mathcal{L}_{u}$ adds a negative term $\frac{1}{T}\sum_{i=1}^{T}\log p(w_{i}|V)$. One may also employ the average bits per character (BPC), a common metric for text compression [27], as the vocabulary-insensitive loss. The only difference lies in the normalization. BPC represents the raw per-character language model loss over the corpus, while our $\mathcal{L}_{u}$ is equivalent to the per-character language model loss normalized by the frequency of each character. Since we employ the same corpus for each model we train, there is not much difference between the two metrics in our case.
@@ 3 Analysis: Why the optimal vocabulary size is bounded by compute
In this section, we present analyses to explain why the optimal vocabulary size is constrained by the computational budget.
@@
3.1 Analysis 1: The perspective of fixed normalized loss
According to Kaplan et al. [30], the FLOPs ($C$) of a Transformer-based language model can be estimated as $C\approx 6ND$, which can be re-written as:
@@ 3.1 Analysis 1: The perspective of fixed normalized loss
3.2 Analysis 2: The perspective of fixed FLOP budget
Given a fixed FLOPs budget, we isolate the FLOPs and investigate how the vocabulary influences the loss. In practice, we set several fixed FLOP budgets. For each budget, we adopt a group of models with similar total parameters and vary vocabulary sizes from 4K to 96K. In Figure 3 (right) we plot the relationship between the loss w.r.t. the vocabulary size. It reveals that the vocabulary corresponding to the lowest point on the loss curve increases as the FLOPs budget increases. This suggests that with more computational resources, LLMs can effectively harness larger vocabularies to reduce loss. However, merely expanding the vocabulary does not always lower the loss. For a fixed FLOP budget, the loss initially decreases with the increase in vocabulary and then starts to rise, indicating that an optimal point exists for the vocabulary. This suggests a trade-off between model complexity and computational constraints, where an overly large vocabulary cannot be utilized efficiently, leading to sub-optimal model performance.
3.3 Analysis 3: The perspective of parameter growing
Non-vocabulary parameters can benefit from increases in both depth and width, allowing for more complex hierarchical representations and broader feature capture. In contrast, vocabulary parameters, associated with word embeddings and language model heads, are generally confined to a single layer, limiting their ability to benefit from increases in the model depth. Their primary avenue for expansion is through increasing the width. This disparity in growth potential between non-vocabulary and vocabulary parameters suggests that to maintain a balanced growth rate, it may be necessary to expand the vocabulary size along with the depth. This would allow the vocabulary parameter part to keep pace with the growth of non-vocabulary parameters.
@@ 4 Estimating the optimal vocabulary size
Figure 4: The training curve envelope of the experiments used in Approach 1 (§4.1) and Approach 3 (§4.3). We train a series of models with the non-vocabulary parameters fixed and vocabulary sizes varying from 4K to 96K.
@@ 4.1 Approach 1: Estimating power laws via IsoFLOPs
Setup
Given a certain $N_{\rm v}$, the embedding dimension $d$ is fixed, thus $N_{\rm v}$ increases as $V$ increases. For all experiments, we uniformly sample the training data from different domains in the SlimPajama dataset [60]. All other hyperparameters are fixed with more details in §A.5.
Fitting
We select data points with the minimum $\mathcal{L}_{u}$ for each FLOP budget, with all runs visualized in Figure 4. These points are the compute-optimal allocation to ($N_{\rm nv}$, $N_{\rm v}$, $H$). Following Kaplan et al. [30] and Hoffmann et al. [26], we hypothesize that the optimal vocabulary parameters $N_{\rm v}$ meet a power law w.r.t. the FLOPs $C$, just like the non-vocabulary parameters and the amount of training data. Specifically, $N_{\rm nv}=k_{1}C^{\alpha_{1}}$,$N_{\rm v}=k_{2}C^{\alpha_{2}}$ and $H=k_{3}C^{\alpha_{3}}$. As model size and training data should be scaled equally for compute-optimal training [26], we set $\alpha_{1}=\alpha_{3}$. As our new attribute $V$ significantly increases the number of possible experimental configurations, we employ interpolation across data points to obtain more configurations at a low cost. The implementation details of the fitting are in §A.5.4.
@@
Results
In Figure 5, we display the fitted power laws: $N_{\rm nv}=0.08*C^{0.50}$, $N_{\rm v}=0.20*C^{0.42}$ and $H=6.42*C^{0.50}$, where $C$ is the FLOPs budget.The low $\rm RMSE$ and high $R^{2}$ values indicate the strength of our fit. Given a certain FLOPs budget, we can utilize the aforementioned relationships to obtain the optimal allocation ($N_{\rm nv}$, $N_{\rm v}$, $H$). We also draw the following conclusions: (1) LLMs are data-hungry. Compared to the non-vocabulary parameters $N_{\rm nv}$, practitioners should allocate more compute to the training data [75, 43]. (2) Vocabulary parameters scale in a power-law relation with FLOPs ($N_{\rm v}\propto C^{0.42}$). As models become more computationally intensive, a larger vocabulary enhances the model’s ability to understand a more diverse array of text, and thus the vocabulary size is critical to scaling. (3) Vocabulary parameters $N_{\rm v}$ should be scaled slower than non-vocabulary parameters $N_{\rm nv}$. This difference can be seen in their power law exponents, i.e. $\gamma=0.42/0.50=0.84<1$. We hypothesize the reason is that: once a sufficiently rich embedding space is present via a large vocabulary, it is more critical to scale non-vocabulary parameters to learn the intricate syntactic and semantic structures of language via Transformer blocks.
@@ 4.2 Approach 2: Derivative-based fast estimation
Usage
Note that the optimal vocabulary size should be determined primarily by the normalized loss $\mathcal{L}_{u}$, rather than by the compute budget FLOPs. However, when the compute allocation is near optimal, the loss exhibits a power-law relationship with respect to the FLOPs budget, as described by the scaling law [30]. This relationship allows us to use FLOPs with compute-optimal allocation as a reliable proxy for observing the scaling behavior of the optimal vocabulary parameters. In practise, we can first determine an empirically optimal vocabulary size in a low-cost setting (e.g., finding the compute-optimal vocabulary parameters on a small model with equivalent FLOPs). Then, we can scale the optimal vocabulary parameters proportionally as we increase the non-vocabulary parameters. Specifically, we obtain a set of derivative-optimal vocabulary parameters $N_{v}$ for different non-vocabulary parameters $N_{nv}$, represented as $\left\{(N_{\rm nv}^{i},N_{\rm v}^{i})|i=1,\cdots,n\right\}$. We then fit the relationship between $N_{\rm nv}$ and $N_{\rm v}$ using the power-law function $N_{\rm v}\propto N_{\rm nv}^{\gamma}$. This results in the scaling equation: $N_{\rm v}/N_{\rm v}^{0}=(N_{\rm nv}/N_{\rm nv}^{0})^{\gamma}$ where $N_{\rm nv}^{0}$ is a relatively small model (e.g., 33M), and $N_{\rm v}^{0}$ is the searched optimal vocabulary parameter with sufficient training characters with the same FLOPs budget. By combining the $\gamma$ value obtained from the derivative and the empirical solution on a small model, we can estimate the optimal vocabulary size for any large model without the need for extensive parameter searches following:

### v3

@@ Abstract
Research on scaling large language models (LLMs) has primarily focused on model parameters and training data size, overlooking the role of vocabulary size. We investigate how vocabulary size impacts LLM scaling laws by training models ranging from 33M to 3B parameters on up to 500B characters with various vocabulary configurations. We propose three complementary approaches for predicting the compute-optimal vocabulary size: IsoFLOPs analysis, derivative estimation, and parametric fit of the loss function. Our approaches converge on the conclusion that the optimal vocabulary size depends on the compute budget, with larger models requiring larger vocabularies. Most LLMs, however, use insufficient vocabulary sizes. For example, we predict that the optimal vocabulary size of Llama2-70B should have been at least 216K, 7 times larger than its vocabulary of 32K. We validate our predictions empirically by training models with 3B parameters across different FLOPs budgets. Adopting our predicted optimal vocabulary size consistently improves downstream performance over commonly used vocabulary sizes. By increasing the vocabulary size from the conventional 32K to 43K, we improve performance on ARC-Challenge from 29.1 to 32.0 with the same 2.3e21 FLOPs. Our work highlights the importance of jointly considering tokenization and model scaling for efficient pre-training. The code and demo are available at https://github.com/sail-sg/scaling-with-vocab and https://hf.co/spaces/sail/scaling-with-vocab-demo.
@@ 1 Introduction
Large language models (LLMs) achieve remarkable performance by pre-training on vast text corpora using massive computational resources [47]. Extensive prior work on LLMs has focused on deriving so-called scaling laws: a set of empirical formulas to predict how model performance scales, mainly as computing floating-point operations (FLOPs), model parameters, and quantity of training data change [30, 26, 66, 2, 44, 58]. These works show that power-law fits can effectively predict language modeling loss and by extension downstream performance [23, 55]. However, these scaling laws usually disregard the impact of the vocabulary size. For example, in Kaplan et al. [30] only non-vocabulary parameters are considered in their predictive formula. This negligence has resulted in substantial variability in the vocabulary size of current LLMs. For instance, Llama2-7B employs a vocabulary size of 32K [70], while Gemma-7B [67] adopts a much larger vocabulary size of 256K despite both having a similar number of total parameters. This variability in vocabulary sizes across LLMs raises the research question: What is the compute-optimal vocabulary size for a LLM?
The vocabulary size affects performance non-trivially. Intuitively, the optimal vocabulary size should neither be too large nor small. A larger vocabulary size improves tokenization fertility, i.e., splitting sentences into fewer tokens, thereby improving the tokenization efficiency. Additionally, a larger vocabulary enables the model to capture a wider range of concept. However, the risk of under-fitting for rare tokens increases with larger vocabulary sizes, especially in the data-constrained regime [44, 72]. Thus, the optimal vocabulary size needs to be determined by taking the training data and the non-vocabulary parameters into account.
@@ 2.1 Scaling law
Following Radford et al. [51], the loss function is typically the language modeling loss when evaluating language models, which can be written as:
@@ 2.2 Scaling law with vocabulary
Attributes    Scaling laws commonly deal with the attributes, model parameters ($N$) and number of training tokens ($D$) [26, 44]. We adapt them for our analysis in the context of vocabulary size. (1) We break down the total model parameters ($N$) into non-vocabulary ($N_{\rm nv}$) and vocabulary parameters ($N_{\rm v}$). To understand the importance of vocabulary parameters, we isolate them from other model parameters, where $N=N_{\rm nv}+N_{\rm v}$. We use $N_{\rm v}=Vd$ to represent both the vocabulary parameters in the output layer  [^1] Vocabulary parameters typically encompass both the word embedding layer and the output layer. In this paper, for clarity and analytical simplicity, we employ $Vd$ rather than $2Vd$ to represent the vocabulary parameters. This choice is predicated on empirical observations: the main computational burden, as measured in FLOPs, is associated with the output layer, but not the word embedding layer.. Notably, to change $N_{\rm v}$ we only vary the vocabulary size $V$ and take the embedding dimension $d$ as given based on $N_{\rm nv}$ empirically, see §A.7.2 for details. This is based on the observation by Kaplan et al. [30] that the performance of models with varying depth-to-width ratios converges to a single trend. We also provide further analysis about why we break down the model parameters from the perspective of parameter growing in §A.6. (2) We measure data not in tokens ($D$) but in training characters ($H$). The number of tokens depends on the vocabulary of the tokenizer. By studying training characters, we can better see how the data volume affects the performance regardless of different vocabulary sizes.
Mapping from training characters ($H$) to tokens ($D$)    As detailed above we measure training data in training characters ($H$). Nonetheless, to connect our findings with existing studies on scaling laws [26, 44], we need to be able to map from $H$ to $D$. This mapping is the tokenizer’s compression ratio which can be computed via $D/H$. The more tokens the tokenizer needs to represent $H$, the larger $D$, and thus it compresses less. We develop a simple function $f(V)$ to estimate this ratio solely from the chosen vocabulary size, $V$. Specifically, we find that a quadratic function on the logarithmic value of $V$ achieves accurate predictions:
@@
Vocabulary-insensitive loss     To fairly assess models that vary in $V$, the commonly used language model loss in Equation 2 is inappropriate. Models trained with larger $V$ naturally have a higher loss, as there are more possibilities in the vocabulary to predict. However, this does not mean that the model is worse. Thus, we need to normalize the loss with respect to the vocabulary size. We reformulate the unigram-normalized metric [54] as a loss function. Suppose we have a $T$-length sequence $w_{1:T}$, we design the unigram-normalized language model loss as:
@@
where $p(w_{i}|V)$ is the frequency of word $w_{i}$ in the tokenized corpus, given the tokenizer with vocabulary size $V$. The loss indicates the improvement in probability that a context-aware language model offers over a unigram model without context, allowing us to assess the language model’s efficacy. Based on theory from prior work [54], the normalized loss $\mathcal{L}_{u}$ remains consistent for a given model with a fixed non-vocabulary component across different vocabulary sizes. The difference of $\mathcal{L}_{u}$ comes from the ability of the language model itself. Compared with $\mathcal{L}$, the value of $\mathcal{L}_{u}$ is much smaller and can be negative as $\mathcal{L}_{u}$ adds a negative term $\frac{1}{T}\sum_{i=1}^{T}\log p(w_{i}|V)$. One may also employ the average bits per character (BPC), a common metric for text compression [27], as the vocabulary-insensitive loss. The only difference lies in the normalization. BPC represents the raw per-character language model loss over the corpus, while our $\mathcal{L}_{u}$ is equivalent to the per-character language model loss normalized by the frequency of each character. In practice, we find that the metric BPC and $\mathcal{L}_{u}$ show a significant positive correlation, which experimentally validated our statement, as detailed in the §A.5.
@@ 3 Analysis: Why the optimal vocabulary size is bounded by compute
Analysis 1: The perspective of fixed normalized loss     According to Kaplan et al. [30], the FLOPs ($C$) of a Transformer model can be estimated as $C\approx 6ND$, which can be re-written as:
@@
Analysis 2: The perspective of fixed FLOP budget     Given a fixed FLOPs budget, we isolate the FLOPs and investigate how the vocabulary influences the loss. For ease, we train models with fixed $N_{nv}$ and different vocabulary sizes for the same steps, and then we use interpolation to predict the loss when FLOPs reaches the budget given the observed FLOPs and loss points. For each budget, we adopt a group of models with similar total parameters and vary vocabulary sizes. In Figure 3 (right) we plot the relationship between the loss w.r.t. the vocabulary size. It reveals that the vocabulary corresponding to the lowest point on the loss curve increases as the FLOPs budget increases. This suggests that with more computational resources, LLMs can harness larger vocabularies to reduce loss. However, merely expanding the vocabulary does not always lower the loss. For a fixed FLOPs budget, the loss initially decreases with the increase in vocabulary and then starts to rise, indicating that an optimal point exists for the vocabulary.
@@ 4.1 Approach 1: Estimating power laws via IsoFLOPs
Figure 4: Training curves of the experiments used in Approach 1 (§4.1) and Approach 3 (§4.3). We train models with the non-vocabulary parameters fixed and vocabulary sizes varying from 4K to 96K.
Setup    Given a certain $N_{\rm v}$, the embedding dimension $d$ is fixed, thus $N_{\rm v}$ increases as $V$ increases. For all experiments, we uniformly sample the training data from different domains in the SlimPajama dataset [61]. All other hyperparameters are fixed with more details in §A.7.
Fitting    We select data points with the minimum $\mathcal{L}_{u}$ for each FLOP budget, with all runs visualized in Figure 4. These points are the compute-optimal allocation to ($N_{\rm nv}$, $N_{\rm v}$, $H$). Following Kaplan et al. [30] and Hoffmann et al. [26], we hypothesize that the optimal vocabulary parameters $N_{\rm v}$ meet a power law w.r.t. the FLOPs $C$, just like the non-vocabulary parameters and the amount of training data. Specifically, $N_{\rm nv}=k_{1}C^{\alpha_{1}}$,$N_{\rm v}=k_{2}C^{\alpha_{2}}$ and $H=k_{3}C^{\alpha_{3}}$. As model size and training data should be scaled equally for compute-optimal training [26], we set $\alpha_{1}=\alpha_{3}$. As our new attribute $V$ significantly increases the number of possible experimental configurations, we employ interpolation across data points to obtain more configurations cheaply. The details of the fitting are in §A.7.4.
@@
Results and Usage     In Figure 5, we display the fitted power laws: $N_{\rm nv}=0.08*C^{0.50}$, $N_{\rm v}=0.20*C^{0.42}$ and $H=6.42*C^{0.50}$, where $C$ is the FLOPs budget.The low $\rm RMSE$ and high $R^{2}$ values indicate the strength of our fit. Given a certain FLOPs budget, we can utilize the aforementioned relationships to obtain the optimal allocation ($N_{\rm nv}$, $N_{\rm v}$, $H$). We also draw the following conclusions: (1) LLMs are data-hungry. Compared to the non-vocabulary parameters $N_{\rm nv}$, practitioners should allocate more compute to the training data [80, 44]. (2) Vocabulary parameters scale in a power-law relation with FLOPs ($N_{\rm v}\propto C^{0.42}$). As models become more computationally intensive, a larger vocabulary enhances the model’s ability to understand a more diverse array of text, and thus the vocabulary size is critical to scaling. (3) Vocabulary parameters $N_{\rm v}$ should be scaled slower than non-vocabulary parameters $N_{\rm nv}$. This difference can be seen in their power law exponents, i.e. $\gamma=0.42/0.50=0.84<1$. We hypothesize the reason is that: once a sufficiently rich embedding space is present via a large vocabulary, it is more critical to scale non-vocabulary parameters to learn the intricate syntactic and semantic structures of language via Transformer blocks.
@@ 4.2 Approach 2: Derivative-based fast estimation
Usage     When the compute allocation is near optimal, the loss exhibits a power-law relationship with respect to the FLOPs budget, as described by the scaling law [30]. This relationship allows us to use FLOPs as a reliable proxy for observing the scaling behavior of the optimal vocabulary parameters. In practice, we can first determine an empirically optimal vocabulary size in a low-cost setting (e.g., finding the compute-optimal vocabulary parameters on a small model). Then, we can scale the optimal vocabulary parameters proportionally based on $\gamma$. Specifically, we obtain a set of derivative-optimal vocabulary parameters $N_{v}$ for different non-vocabulary parameters $N_{nv}$, represented as $\left\{(N_{\rm nv}^{i},N_{\rm v}^{i})|i=1,\cdots,n\right\}$. We then fit the relationship between $N_{\rm nv}$ and $N_{\rm v}$ using the power-law function $N_{\rm v}\propto N_{\rm nv}^{\gamma}$. This results in the scaling equation: $N_{\rm v}/N_{\rm v}^{0}=(N_{\rm nv}/N_{\rm nv}^{0})^{\gamma}$ where $N_{\rm nv}^{0}$ is a small model (e.g., 33M), and $N_{\rm v}^{0}$ is the searched optimal vocabulary parameter. By combining the $\gamma$ from the derivative and the empirical solution on a small model, we can estimate the optimal vocabulary by: